OpenStudy (anonymous):

A balloon is a sphere with a radius of 5.0 m. The force of air against the walls of the balloon is 45 N. What is the air pressure inside the balloon? (surface area = 12.56 x radius^2; 1 Pa = 1 N/m^2) A. 1.4 x 10^3 Pa B. 3.6 x 10^2 Pa C. 1.4 x 10^-1 Pa D. 3.6 x 10^-2 Pa ***Idk how to solve this... pls explain?? thx :)

jimthompson5910 (jim_thompson5910):

Do you know the surface area of a sphere?

OpenStudy (anonymous):

umm is it 4(pi)(r^2) ??? idk lol

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so the surface area of the balloon is ???

OpenStudy (anonymous):

4(3.14)(5^2) =4(3.14)(25) =100(3.14) =314 ??

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

that's the surface area in square meters

OpenStudy (anonymous):

mmk :)

OpenStudy (anonymous):

what do i do next? :O

jimthompson5910 (jim_thompson5910):

find the pressure in pascals 1 Pa = 1 N/m^2

jimthompson5910 (jim_thompson5910):

so effectively, you're dividing the force in newtons by the surface area in m^2

OpenStudy (anonymous):

oh okay, what do i plug in tho??

OpenStudy (anonymous):

they gave me 45 N.... idk what else?? :/

jimthompson5910 (jim_thompson5910):

you just found the m^2, it's 314 so 45/314 = ???

OpenStudy (anonymous):

ohhh 45/314=0.143312102

OpenStudy (anonymous):

so u turn it into 1.4 ??

jimthompson5910 (jim_thompson5910):

0.14 = 1.4 x 10^-1, so it's C

jimthompson5910 (jim_thompson5910):

and you were correct

OpenStudy (anonymous):

yay!! :) thank youuu :D

OpenStudy (anonymous):

answer C it is then!! :D thx :)